OpenStudy (anonymous):

Solve by elimination: 4x-y=-5 -2x+3y=10

OpenStudy (thatonegirl_):

To solve by elimination, you need to have one equation with a negative coefficient (and same variable) along with the other equation but with the positive coefficient (and same variable). To get the same coefficients, you need to multiply the entire equation by a number so it is equal to the other equations coefficient.

OpenStudy (anonymous):

Okay so if I multiply the top one by 3 would that work?

OpenStudy (thatonegirl_):

So in the first equation there is a 4x, and the second one there is a -2, so you could multiply (-2x+3y=10)*2 so that you get -4x-12y=40.

OpenStudy (thatonegirl_):

or ya u can do the y too if u want u would multiply it all by 3 like u said

OpenStudy (thatonegirl_):

Do you know what to do after that?

OpenStudy (anonymous):

but… when I do that i end up getting -6x=-15 and that is really weird.

OpenStudy (anonymous):

Oh… never mind I figured out what I was doing wrong! Thanks for you help!

OpenStudy (thatonegirl_):

yea no prob! Did u end up getting x=-1/2?

OpenStudy (anonymous):

Yes.

OpenStudy (thatonegirl_):

okay cool, then u just plug it into either of the equations to find y :)

OpenStudy (anonymous):

Okay, thanks again!

OpenStudy (thatonegirl_):

You're welcome!